Show that f(x y)=xy^(2) satisfies a Lipschite condition on any rectangle a<=x<b and c<=y<=d.


1
Expert's answer
2021-07-14T12:07:43-0400

A function f(x, y) is said to satisfy Lipschitz condition on a domain "D\sube\R^2" , if there exists "L>0" such that "|f(x, y_1)-f(x, y_2)|\leq L|y_1-y_2|" for all "(x, y_1), (x, y_2)\in D."

On closed rectangle "[a, b]\times[c, d]," the partial derivative "f_y" is continuous and hence bounded and hence "f" satisfies Lipschitz condition. 


Note that "|x|\leq max(|a|, |b|)=A" and "|y|\leq max(|c|, |d|)=C."



"|\dfrac{f(x, y_1)-f(x, y_2)}{y_1-y_2}|=|\dfrac{xy_1^2-xy_2^2}{y_1-y_2}|"

"=|x(y_1+y_2)|\leq|x||y_1+y_2|\leq2A\cdot C"

is bound.


The function "f(x, y)=xy^2" satisfies Lipschitz condition on any rectangle "[a, b]\times[c, d]."